# Neural Diffusion Models

### v2

@@ B.4 DiffEnc
B.5 Discussion
In theory, we can view the diffusion model as a hierarchical VAE. From this perspective, the conventional diffusion model can be seen as a VAE with a fixed variational distribution. In contrast NDMs with learnable transformations have the capability to make this distribution learnable. Consequently, it effectively reduces the gap between the log likelihood and the variational bound.
In practical terms, when dealing with reverse processes parameterised by neural networks, which inherently possess their own biases, the learnable $F_{\varphi}$ function tends to acquire a transformation that facilitates the fitting of the reverse process to the forward process. Therefore, as soon as there are infinitely many pairs of forward and reverse processes with same joint distributions of latent variables, it’s hard to say what $F_{\varphi}$ should be in general case. It may learn anything that fits to the reverse process. However, besides better likelihood estimation, NDMs with learnable transformations gives us additional flexibility in how we can define the reverse process.
@@ Appendix C Implementation details
All our experiments were conducted using synthetic 2D datasets and image datasets: MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), downsampled ImageNet (Deng et al., 2009) and CelebA-HQ-256 (Karras et al., 2017). For CIFAR-10 and ImageNet datasets we applied center cropping and resizing. For synthetic data, we employed a 5-layer MLP with 512 neurons in each layer, while for the images, we utilized the U-Net architecture from Dhariwal & Nichol (2021). In our experiments both the DDPM and NDM approaches were trained on identical architectures, with the same hyper-parameters and for the same number of epochs. The hyper-parameters are presented in Table 6. In experiment where we report results for the continuous time models we use importance sampling of time (Song et al., 2021) instead of uniform sampling.
@@ D.2 Additional samples
Figure 4 provides a comparison between DDPM and NDM on a synthetic 2D data distribution. For this experiment, both models utilize $T=10$ discrete time steps. From Figure 3(c), it is evident that NDM learns to transform the data distribution. Additionally, after injecting noise (Figures 3(a) and 3(d)), the distributions of samples $\mathbf{z}_{t}$ show minimal differences between DDPM and NDM. However, when examining the predictions of data points $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (Figures 3(b) and 3(e)), NDM produces predictions that more closely resemble the true data distribution compared to DDPM.
@@ E.1 Restricted reverse process
Next, we may switch to a stochastic differential equation (SDE) according to Song et al. (2020b):
@@ E.3 Results and discussion
Figures 2(a) and 2(b) illustrate trajectories learned by DDPM and NDM with learnable $F_{\varphi}$ and restricted reverse process. As expected, DDPM learns curved trajectories predetermined by fixed forward process. At the same time NDM effectively learns dynamic OT. It worths noting that DDPM with the restricted reverse process is by design not able to learn the data distribution, since it’s impossible to match the fixed forward process (with curved trajectories) with the reverse process (with straight trajectories).

### v3

@@ Appendix C Implementation details
All our experiments were conducted using synthetic 2D datasets and image datasets: MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) and CelebA-HQ-256 (Karras et al., 2017). For CIFAR-10 and ImageNet datasets we applied center cropping and resizing. For synthetic data, we employed a 5-layer MLP with 512 neurons in each layer, while for the images, we utilized the U-Net architecture from Dhariwal & Nichol (2021). In our experiments both the DDPM and NDM approaches were trained on identical architectures, with the same hyper-parameters and for the same number of epochs. The hyper-parameters are presented in Table 6. In experiment where we report results for the continuous time models we use importance sampling of time (Song et al., 2021b) instead of uniform sampling.
@@ D.2 Additional samples
Figure 4 provides a comparison between DDPM and NDM on a synthetic 2D data distribution. For this experiment, both models utilize $T=10$ discrete time steps. From Figure 4(c), it is evident that NDM learns to transform the data distribution. Additionally, after injecting noise (Figures 4(a) and 4(d)), the distributions of samples $\mathbf{z}_{t}$ show minimal differences between DDPM and NDM. However, when examining the predictions of data points $\hat{\mathbf{x}}_{\theta}(\mathbf{z}_{t},t)$ (Figures 4(b) and 4(e)), NDM produces predictions that more closely resemble the true data distribution compared to DDPM.
@@ E.1 Restricted reverse process
Next, we may switch to a stochastic differential equation (SDE) according to Song et al. (2021c):
@@ E.3 Results and discussion
Figures 3(a) and 3(b) illustrate trajectories learned by DDPM and NDM with learnable $F_{\varphi}$ and restricted reverse process. As expected, DDPM learns curved trajectories predetermined by fixed forward process. At the same time NDM effectively learns dynamic OT. It worths noting that DDPM with the restricted reverse process is by design not able to learn the data distribution, since it’s impossible to match the fixed forward process (with curved trajectories) with the reverse process (with straight trajectories).